# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ Abstract
Patterns are words with terminals and variables. The language of a pattern is the set of words obtained by uniformly substituting all variables with words that contain only terminals. In their original definition, patterns only allow for multiple distinct occurrences of some variables to be related by the equality relation, represented by using the same variable multiple times. In an extended notion, called relational patterns and relational pattern languages, variables may be related by arbitrary other relations. We extend the ongoing investigation of the main decision problems for patterns (namely, the membership problem, the inclusion problem, and the equivalence problem) to relational pattern languages under a wide range of individual relations. It is shown show that - even for many much simpler or less restrictive relations - the complexity and (un)decidability characteristics of these problems do not change compared to the classical case where variables are related only by equality.
@@ 1 Introduction
Originally, pattern languages as defined and introduced by Angluin [Angluin1980] only considered words that are obtained by what is now called non-erasing substitutions, i.e., substitutions where each variable is substituted by a word of at least length one. Hence, these languages are now also called non-erasing (NE) pattern languages. Shinohara [Shinohara1983] extended this notion by allowing for the empty word to be used to substitute a variable, introducing so called extended/erasing (E) pattern languages.
@@
As patterns form a natural and compact device for the definition of formal languages and due to their practical and simple definition, they occur in numerous areas in computer science and discrete mathematics. Among others, these include for example the areas of unavoidable patterns in words [Jiang1994, lothaire1997], word equations [lothaire1997], algorithmic learning theory [Angluin1980, FERNAU201844, Shinohara1995], database theory [FreydenbergerP21, SchmidSchweikardtPODS2022], or the theory of extended regular expressions with back references[FREYDENBERGER20191]. Furthermore, many practical areas such as machine learning, database systems, or bio-informatics deal with the question of identifying or describing patterns in (sets of) strings [Mousawi2024], for example in the analysis of protein data [Arikawa1993], in pattern matching [10.1007/978-3-642-03784-9_29], or the design of algorithms for program synthesis [10.1145/4472.4476].
@@
Three main decision problems emerge when it comes to the investigation of patterns and pattern languages. These are the general membership problem and its variations [gawrychowski_et_al:LIPIcs.MFCS.2021.48, Manea2022, Fleischmann2023], the inclusion problem, and the equivalence problem, all of which are considered in the non-erasing (NE) as well as erasing (E) cases. The membership problem determines whether a word belongs to a pattern’s language. It has been shown to be generally NP-complete in the non-erasing as well as erasing cases [Angluin1980, Jiang1994]. The inclusion problem checks whether the language of a pattern is included as a subset in another pattern’s language. It has been shown to be generally undecidable by Jiang et al. [Jiang1995]. Their result has been improved by Freydenberger and Reidenbach in [Freydenberger2010] as well as by Bremer and Freydenberger in [BREMER201215], where it has been shown to be undecidable for all alphabets $\Sigma$ of fixed size $|\Sigma|\geq 2$, for erasing and non-erasing pattern languages. The equivalence problem determines whether the languages of two given patterns are equal. In Angluin’s original work [Angluin1980], it has been shown to be trivially decidable for all NE pattern languages. For erasing pattern languages, however, the general decidability of this problem has been the most significant open problem in the field of decision problems for patterns [Jiang1995, Reidenbach2004-1, Reidenbach2004-2, Ohlebusch1996, Reidenbach2007]. In the special case of terminal-free patterns (containing only variables), the membership problem remains unchanged in both erasing and non-erasing cases, as does the non-erasing equivalence problem. However, inclusion for terminal-free E-pattern languages is NP-complete, as it reduces to checking for a morphism between patterns [Jiang1995, DBLP:journals/ipl/EhrenfeuchtR79a], while equivalence is decidable. In contrast, Saarela [saarela:LIPIcs.ICALP.2020.140] showed that inclusion for terminal-free NE pattern languages is undecidable over a binary alphabet.
@@
Various extensions to the notion of patterns and pattern languages have been introduced over time either to get closer to an answer for the remaining open problems or to obtain additional expressibility that is usable in some practical context. Some examples are the bounded scope coincidence degree, patterns with bounded treewidth, $k$-local patterns, or strongly-nested patterns (see [Day2018] and references therein). Koshiba [Koshiba1995] introduced so called typed patterns where the substitutions of variables can be restricted by arbitrary recursive languages, so called types. This has been recently extended by Geilke and Zilles in [Geilke2011] to the notion of relational patterns and relational pattern languages. Here, other relations than equality between variables may be used to restrict the valid substitution of patterns.
@@
This model enhances expressiveness, includes all previous models, but also includes arbitrary other constructions that involve relations between variables in a pattern. Consider, for example, the pattern $\beta=x_{1}\mathtt{c}\mathtt{c}x_{2}$. Now let’s say that the variables $x_{1}$ and $x_{2}$ should be the substituted by the reversal of each other, i.e., $(h(x_{1}),h(x_{2}))\in r_{rev}$, where $h$ is a substitution and $r_{rev}$ is the reversal relation defined by $r_{rev}:=\{\ (w,w^{R})\mid w\in\Sigma^{*}\ \}$ (where $w^{R}$ is defined to be the reversal of a word $w$). Then, e.g., a substitution $h$ with $h(x_{1})=\mathtt{a}\mathtt{b}$ and $h(x_{2})=\mathtt{b}\mathtt{a}$ is said to be valid regarding the pattern $\beta$, we get $h(\beta)=\mathtt{a}\mathtt{b}\mathtt{c}\mathtt{c}\mathtt{b}\mathtt{a}$, and we say that $h(\beta)$ is in the relational pattern language of $\beta$ under the relational constraints containing the pair that relates $x_{1}$ and $x_{2}$ by $r_{rev}$. If we considered a relation $h^{\prime}$, however, in which $h^{\prime}(x_{1})=\mathtt{a}\mathtt{b}$ and also $h^{\prime}(x_{2})=\mathtt{a}\mathtt{b}$, then clearly $h^{\prime}(x_{1})\neq h^{\prime}(x_{2})^{R}$ and therefore $h^{\prime}$ is not a valid substitution in this case, resulting in $h^{\prime}(\beta)$ not being in that relational pattern language of $\beta$.
@@
Geilke and Zilles [Geilke2011] show some properties of relational pattern languages in general. Holte et al. [pmlr-v167-holte22a] investigated relational pattern languages under the equal length or reversal relations, respectively. They show that inclusion, by that also equivalence, under equal length is decidable in polynomial time for ternary alphabets. In a very recent paper, Mousawi and Zilles [Mousawi2024] extended this from the perspective of learning theory, introducing the notion of positive characteristic sets. Their results carry over and it is shown that inclusion, therefore also equivalence, in the erasing and non-erasing cases under equal length is decidable in polynomial time for binary alphabets. Also, for the reversal relation, their result implies that in the non-erasing case, equivalence is decidable. Additionally, also in two recent papers, in the cases of two special kinds of relations/constraints, namely regular constraints [Nowotka2024] or length constraints[Nowotka2025], Nowotka and Wiedenhöft have shown, among other results, that the still open equivalence problem for erasing pattern languages actually becomes undecidable if these constraints are allowed in addition to equality between variables. This sets a first upper bound of constraints necessary to obtain undecidability for that problem. In comparison to many other relations, however, these constraints appear to be very strong, hence leaving the question of what happens in cases of weaker relations.
@@
In total, we see that the field of relational pattern languages remains largely unexplored, leaving room to study various other or weaker relations restricting variable substitutions. This paper considers equality ($r_{=}$), equal length ($r_{|w|}$), subsequence ($r_{ssq}$), abelian equivalence ($r_{ab}$), alphabet permutation ($r_{perm}$), reversal ($r_{rev}$), (nonempty) commutation ($r_{com^{+}}$, $r_{com^{*}}$), and the star relation ($r_{*}$). We study the membership, inclusion, and equivalence problems under these relations, providing a comprehensive overview of many cases (Table 3) and showing that hardness and (un)decidability persist even for weaker relations than equality. In particular, it can be shown that the membership problem is NP-complete for all considered relations in all cases. Regarding the inclusion problem, it is shown that undecidability persists even for the relations $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ in the erasing and non-erasing cases. Finally, in the case of the equivalence problem, we show decidability in polynomial time in the non-erasing case for a class of relations with specific properties, containing $r_{ab}$ and $r_{com^{+}}$.
@@
Section 2 covers all the necessary notation and models needed to follow the results of the paper. In Section 3, we present the results for the membership problem. This is followed in Section 4 with an overview of the results for the inclusion problem. After that, in Section 5, we briefly discuss the results regarding the equivalence problem. Finally, in Section 6, we summarize the paper and provide a brief outlook with regards to open questions. Due to space constraints, some proofs were moved into the appendix. The corresponding statements are marked with (✂).
@@ 2 Preliminaries
Let $\mathbb{N}$ denote the natural numbers $\{1,2,3,\dots\}$ and let $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $n,m\in\mathbb{N}$ set $m,n]:=\{k\in\mathbb{N}\mid m\leq k\leq n\}$. Denote $[n]:=[1,n]$ and $n]_{0}:=[0,n]$. An alphabet $\Sigma$ is a non-empty finite set whose elements are called letters. A word is a finite sequence of letters from $\Sigma$. Let $\Sigma^{*}$ be the set of all finite words over $\Sigma$ and set $\Sigma^{+}:=\Sigma^{*}\setminus\{\varepsilon\}$. We call the number of letters in a word $w\in\Sigma^{*}$ length of $w$, denoted by $|w|$. Therefore, we have $|\varepsilon|=0$. Denote $\Sigma^{k}:=\{w\in\Sigma^{*}\mid|w|=k\}$. For $w\in\Sigma^{*}$, let $w[i]$ denote $w$’s $i^{th}$ letter for all $i\in[|w|]$. For compactness reasons, we denote $w[i]\cdots w[j]$ by $w[i\cdots j]$ for all $i,j\in[|w|]$ with $i<j$. Set $\operatorname{alph}(w):=\{\mathtt{a}\in\Sigma\mid\exists i\in[|w|]:w[i]=\mathtt{a}\}$ as $w$’s alphabet.
@@
Let $X:=\{x_{1},x_{2},$...\} be a countable set of variables such that $\Sigma\cap X=\emptyset$. A pattern is then a non-empty, finite word over $\Sigma\cup X$. The set of all patterns over $\Sigma\cup X$ is denoted by $Pat_{\Sigma}$. For example, $x_{1}\mathtt{a}x_{2}\mathtt{b}\mathtt{a}x_{2}x_{3}$ is a pattern over $\Sigma=\{\mathtt{a},\mathtt{b}\}$ with $x_{1},x_{2},x_{3}\in X$. For a pattern $\alpha\in Pat_{\Sigma}$, let $\operatorname{var}(\alpha):=\{\ x\in X\ |\ |p|_{x}\geq 1\ \}$ denote the set of variables occurring in $p$. A substitution of $\alpha$ is a morphism $h:(\Sigma\cup X)^{*}\to\Sigma^{*}$ such that $h(\mathtt{a})=\mathtt{a}$ for all $\mathtt{a}\in\Sigma$ and $h(x)\in\Sigma^{*}$ for all $x\in X$. If we have $h(x)\neq\varepsilon$ for all $x\in\operatorname{var}(p)$, we call $h$ a non-erasing substitution for $\alpha$. Otherwise $h$ is an erasing substitution for $\alpha$. The set of all substitutions w.r.t. $\Sigma$ is denoted by $H_{\Sigma}$. If $\Sigma$ is clear from the context, we may write just $H$. Given a pattern $\alpha\in Pat_{\Sigma}$, it’s erasing pattern language $L_{E}(\alpha)$ and its non-erasing pattern language $L_{NE}(\alpha)$ are defined respectively by
$L_{E}(\alpha)$ $:=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha)\},\text{ and }$
$L_{NE}(\alpha)$ $:=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\}.$
@@
Let $RegPat_{\Sigma}:=\{\alpha\in Pat_{\Sigma}\mid x_{i}\in\operatorname{alph}(\alpha)\cap X,|\alpha|_{x_{i}}=1\}$ denote the set of all regular patterns, i.e., the set of all patterns where each variable occurs only once. Let $R$ be some set of relations and denote for each $n\in\mathbb{N}$ by $R_{n}\subset R$ the subset of $n$-ary relations in $R$. Define $Rel_{R}:=\{(r,x_{i_{1}},...,x_{i_{n}})\mid r\in R_{n},x_{i_{1}},...,x_{i_{n}}\in X,i_{1},...i_{n}\in\mathbb{N},n\in\mathbb{N}\}$ to be the set of all combinations to relate variables in $X$ over the set of relations $R$. Then, a pair $(\alpha,r_{\alpha})\in RegPat_{\Sigma}\times Rel_{R}$ is a relational pattern over the terminal alphabet $\Sigma$ and the set of relations $R$. We denote the set of all relational patterns over $\Sigma$ and $R$ by $\mathtt{RelPat}_{\Sigma,R}$. For some $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $h\in H$, we say that $h$ is a $r_{\alpha}$-valid substitution if for all $(r^{\prime},x_{i_{1}},...,x_{i_{n}})\in r_{\alpha}$ we have $(h(x_{i_{1}}),...,h(x_{i_{n}}))\in r^{\prime}$, i.e., all variables are substituted such that the resulting terminal words are related to each other regarding the relation $r^{\prime}$. The set of all $r_{\alpha}$-valid substitutions w.r.t. $\Sigma$ is denoted by $H_{\Sigma,r_{\alpha}}$. If $\Sigma$ and $r_{\alpha}$ are clear from the context, we may also just write $H_{r_{\alpha}}$ or $H$. We extend the notion of pattern languages by the following. For any $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ we denote by
$L_{E}(\alpha,r_{\alpha}):=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha),h\text{ is }r_{\alpha}\text{-valid}\ \}$
the erasing relational pattern language of $(\alpha,r_{\alpha})$ and by
$L_{NE}(\alpha,r_{\alpha}):=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha),h\text{ is }r_{\alpha}\text{-valid}\ \}$
the non-erasing relational pattern language of $(\alpha,r_{\alpha})$.
@@
In this paper, we primarily investigate the considered decision problems for relational pattern languages for binary relations, i.e., $|R|=1$ and if $R=\{\tilde{r}\}$, then $\tilde{r}$ is 2-ary/binary. In particular, we are interested in the following selection of specific customary binary relations. Let $\mathcal{R}$ be the set containing all of the following nine relations.

### v2

@@ Abstract
Patterns are words with terminals and variables. The language of a pattern is the set of words obtained by uniformly substituting all variables with words that contain only terminals. In their original definition, patterns only allow for multiple distinct occurrences of some variables to be related by the equality relation, represented by using the same variable multiple times. In an extended notion, called relational patterns and relational pattern languages, variables may be related by arbitrary other relations. We extend the ongoing investigation of the main decision problems for patterns (namely, the equivalence problem, the inclusion problem, and the membership problem) to relational pattern languages under a wide range of relevant individual relations, providing a comprehensive foundation in all three research directions.
@@ 1 Introduction
Originally, pattern languages as defined and introduced by Angluin [3] only considered words that are obtained by what is now called non-erasing substitutions, i.e., substitutions where each variable is substituted by a word of at least length one. Hence, these languages are now also called non-erasing (NE) pattern languages. Shinohara [42] extended this notion by allowing for the empty word to be used to substitute a variable, introducing so-called extended/erasing (E) pattern languages.
@@
As patterns form a natural and compact device for the definition of formal languages and due to their practical and simple definition, they occur in numerous areas in computer science and discrete mathematics. Among others, these include for example the areas of unavoidable patterns in words [22, 26], word equations [26], algorithmic learning theory [3, 9, 43], database theory [14, 41], or the theory of extended regular expressions with back references[15]. Furthermore, many practical areas such as machine learning, database systems, or bio-informatics deal with the question of identifying or describing patterns in (sets of) strings [30], for example in the analysis of protein data [4], in pattern matching [6], or the design of algorithms for program synthesis [32].
@@
Three main decision problems emerge when it comes to the investigation of patterns and pattern languages. These are the general equivalence problem, the inclusion problem, and the membership problem and its variations [16, 17, 12], all of which are considered in the non-erasing (NE) as well as erasing (E) cases. The equivalence problem determines whether the languages of two given patterns are equal. In Angluin’s original work [3], it has been shown to be trivially decidable for all NE pattern languages. For erasing pattern languages, however, the general decidability of this problem has been the most significant open problem in the field of decision problems for patterns [23, 38, 37, 35, 39]. The inclusion problem checks whether the language of a pattern is included as a subset in another pattern’s language. It has been shown to be generally undecidable by Jiang et al. [23]. Their result has been improved by Freydenberger and Reidenbach in [13] as well as by Bremer and Freydenberger in [5], where it has been shown to be undecidable for all alphabets $\Sigma$ of fixed size $|\Sigma|\geq 2$, for erasing and non-erasing pattern languages. The membership problem determines whether a word belongs to a pattern’s language. It has been shown to be generally NP-complete in the non-erasing as well as erasing cases [3, 22] and has been investigated extensively in various communities [2, 36, 11, 10, 20, 28, 27]. In the special case of terminal-free patterns (containing only variables), the membership problem remains hard in both erasing and non-erasing cases, as does the non-erasing equivalence problem. However, inclusion for terminal-free E-pattern languages is NP-complete, as it reduces to checking for a morphism between patterns [23, 8], resulting also in decidability of equivalence. In contrast, Saarela [40] showed that inclusion for terminal-free NE pattern languages is undecidable over a binary alphabet.
@@
Various extensions to the notion of pattern languages have been introduced over time either to get closer to an answer for the remaining open problems or to obtain additional expressibility that is usable in some practical context. Some examples include the bounded scope coincidence degree, patterns with bounded treewidth, $k$-local patterns, or strongly-nested patterns (see [7] and references therein). Koshiba [25] introduced so-called typed patterns where the substitutions of variables can be restricted by arbitrary recursive languages, so-called types. This has been recently extended by Geilke and Zilles in [18] to the notion of relational patterns and relational pattern languages. Here, relations other than equality may be used to restrict the valid substitution of patterns.
@@
This model enhances expressiveness, includes all previous models, but also includes arbitrary other constructions that involve relations between variables in a pattern. Consider, for example, the pattern $\beta=x_{1}\mathtt{c}\mathtt{c}x_{2}$. Assume $x_{1}$ and $x_{2}$ should be substituted by the reversal of each other, i.e., $(h(x_{1}),h(x_{2}))\in r_{rev}$, where $h$ is a substitution and $r_{rev}$ is the reversal relation defined by $r_{rev}:=\{\ (w,w^{R})\mid w\in\Sigma^{*}\ \}$ (where $w^{R}$ is the reversal of a word $w$). Then, a substitution $h$ with $h(x_{1})=\mathtt{a}\mathtt{b}$ and $h(x_{2})=\mathtt{b}\mathtt{a}$ is called valid regarding the pattern $\beta$, we get $h(\beta)=\mathtt{a}\mathtt{b}\mathtt{c}\mathtt{c}\mathtt{b}\mathtt{a}$, and we say that $h(\beta)$ is in the relational pattern language of $\beta$ under the relational constraints that relate $x_{1}$ and $x_{2}$ by $r_{rev}$. If we considered a substitution $h^{\prime}$, however, in which $h^{\prime}(x_{1})=\mathtt{a}\mathtt{b}$ and also $h^{\prime}(x_{2})=\mathtt{a}\mathtt{b}$, then clearly $h^{\prime}(x_{1})\neq h^{\prime}(x_{2})^{R}$ and therefore $h^{\prime}$ is not a valid substitution in this case, resulting in $h^{\prime}(\beta)$ not being in that relational pattern language of $\beta$.
@@
Geilke and Zilles [18] show some properties of relational pattern languages in general. Holte et al. [19] investigated relational pattern languages under the equal length or reversal relations, respectively. They show that inclusion, and thus also equivalence, under equal length is decidable in polynomial time for ternary alphabets. Recently, Mousawi and Zilles [30] extended this from the perspective of learning theory, introducing the notion of positive characteristic sets. Their results carry over and it is shown that inclusion, therefore also equivalence, in the erasing and non-erasing cases under equal length is decidable in polynomial time for binary alphabets. Also, for the reversal relation, their result implies that in the non-erasing case, equivalence is decidable. Additionally, in the cases of two special kinds of relations/constraints, namely regular constraints [33] or length constraints[34], it was recently shown, among other results, that the still open equivalence problem for erasing pattern languages actually becomes undecidable if these constraints are allowed in addition to equality between variables. This sets a first upper bound of constraints necessary to obtain undecidability for that problem. In comparison to other relations, however, these constraints appear to be very strong, leaving the question of what happens in cases of weaker relations.
@@
Hence, the field of relational pattern languages remains largely unexplored, leaving room to study various other relations. This paper considers equality ($r_{=}$), equal length ($r_{|w|}$), subsequence ($r_{ssq}$), abelian equivalence ($r_{ab}$), alphabet permutation ($r_{perm}$), reversal ($r_{rev}$), (nonempty) commutation ($r_{com^{+}}$, $r_{com^{*}}$), the star relation ($r_{*}$), and the common (non)emptiness relation ($r_{\varepsilon_{=}}$). We study the equivalence, inclusion, and membership problems under these relations, providing a comprehensive overview of many cases (Table 1), with a primary focus on general patterns that may contain terminals. We show that in some cases interesting decidability results emerge. In many other cases, however, it is shown that hardness and (un)decidability persist even for structurally weaker relations than equality. In particular, it can be shown that, first, the equivalence problem in the non-erasing cases is decidable in polynomial time for $r_{\varepsilon_{=}}$, $r_{ab}$ and $r_{com^{+}}$ and that equivalence of pattern languages over the relations $_{ssq}, rr$$_{rev}$, and $r_{*}$ emit strong structural restrictions. For erasing cases, i.e. the most important open question for $r_{=}$, that problem turns out to be decidable for $r_{\varepsilon_{=}}$. On the other hand, it is shown that this problem must be at least as hard as for $r_{=}$ for the relations $r_{rev}$, $r_{ssq}$, and $r_{*}$. Regarding the inclusion problem, we obtain decidability in the case of $r_{\varepsilon_{=}}$ while also observing that undecidability persists even for the relations $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ in the erasing and non-erasing cases. Finally, it can be shown that the membership problem is NP-complete for all considered relations in all cases, except for $r_{\varepsilon_{=}}$ in the non-erasing case.
@@
Section 2 covers all the necessary notation and models needed to follow the results of the paper. In Section 3, we discuss the equivalence problem. This is followed in Section 4 with a discussion of the inclusion problem. After that, in Section 5, we briefly discuss the results for the membership problem. Finally, in Section 6, we provide a final discussion of the established results and worthwhile research direction that could follow from the open cases. Due to space constraints, some proofs were moved into the appendix, marked with (✂) and a reference to the respective subsection.
$r_{=}$ $r_{|w|}$ $r_{ssq}$ $r_{ab}$ $_{perm} rr$$_{rev}$ $r_{com^{*}}$ $_{com^{+}} rr_{*}$ $r_{\varepsilon_{=}}$
 E ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC NPC
 E ($\subseteq$)  UD  P UD UD ? UD ? ? UD PS
 E (==) ?  P ? ? ? ? ? ? ? PS
 NE ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC P
 NE ($\subseteq$)  UD  P UD UD ? UD ? ? UD P
 NE (==)  P  P ? P ?  D P P ? P
Table 1: Current state of decision problems for relational pattern languages for various constraints (PS = PSPACE, NPC = NP-Complete, P = Poly. time decidable, UD = Undecidable, D = Decidable, E = Erasing, NE = Non-Erasing, ($\in$) membership problem, ($\subseteq$) inclusion problem, (==) equivalence problem). Red text marks the results covered in this paper.
@@ 2 Preliminaries
Let $\mathbb{N}$ denote the natural numbers $\{1,2,3,\dots\}$ and let $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $n,m\in\mathbb{N}$ set $m,n]:=\{k\in\mathbb{N}\mid m\leq k\leq n\}$. Denote $[n]:=[1,n]$ and $n]_{0}:=[0,n]$. An alphabet $\Sigma$ is a non-empty finite set whose elements are called letters. A word is a finite sequence of letters from $\Sigma$. Let $\Sigma^{*}$ be the set of all finite words over $\Sigma$ and set $\Sigma^{+}:=\Sigma^{*}\setminus\{\varepsilon\}$. We call the number of letters in a word $w\in\Sigma^{*}$ length of $w$, denoted by $|w|$. Therefore, we have $|\varepsilon|=0$. Denote $\Sigma^{k}:=\{w\in\Sigma^{*}\mid|w|=k\}$. For $w\in\Sigma^{*}$, let $w[i]$ denote $w$’s $i^{th}$ letter for all $i\in[|w|]$. For compactness reasons, we denote $w[i]\cdots w[j]$ by $w[i\cdots j]$ for all $i,j\in[|w|]$ with $i<j$. Denote by $|w|_{\mathtt{a}}=|\{\ i\in[|w|]\mid w[i]=\mathtt{a}\}|$ the number of times the letter $\mathtt{a}$ occurs in $w$. Set $\operatorname{alph}(w):=\{\mathtt{a}\in\Sigma\mid\exists i\in[|w|]:w[i]=\mathtt{a}\}$ as $w$’s alphabet. A subsequence $u$ of a word $w$ is a word obtained by deleting arbitrary letters of $w$. Denote by $\operatorname{SubSeq}(w)$ the set of subsequences of $w$. Two words $u,v\in\Sigma^{*}$ are abelian equivalent, denoted by $u\equiv_{ab}v$, iff $|u|_{\mathtt{a}}=|v|_{\mathtt{a}}$ for all $\mathtt{a}\in\operatorname{alph}(u)\cup\operatorname{alph}(v)$.
@@
Let $X:=\{x_{1},x_{2},$...\} be a countable set of variables such that $\Sigma\cap X=\emptyset$. A pattern is then a non-empty, finite word over $\Sigma\cup X$. The set of all patterns over $\Sigma\cup X$ is denoted by $Pat_{\Sigma}$. For example, $x_{1}\mathtt{a}x_{2}\mathtt{b}\mathtt{a}x_{2}x_{3}$ is a pattern over $\Sigma=\{\mathtt{a},\mathtt{b}\}$ with $x_{1},x_{2},x_{3}\in X$. For a pattern $\alpha\in Pat_{\Sigma}$, let $\operatorname{var}(\alpha):=\{\ x\in X\ |\ |p|_{x}\geq 1\ \}$ denote the set of variables occurring in $p$. A substitution of $\alpha$ is a morphism $h:(\Sigma\cup X)^{*}\to\Sigma^{*}$ such that $h(\mathtt{a})=\mathtt{a}$ for all $\mathtt{a}\in\Sigma$ and $h(x)\in\Sigma^{*}$ for all $x\in X$. If we have $h(x)\neq\varepsilon$ for all $x\in\operatorname{var}(p)$, we call $h$ a non-erasing substitution for $\alpha$. Otherwise $h$ is an erasing substitution for $\alpha$. The set of all substitutions w.r.t. $\Sigma$ is denoted by $H_{\Sigma}$. If $\Sigma$ is clear from the context, we may write just $H$. Given a pattern $\alpha\in Pat_{\Sigma}$, its erasing pattern language $L_{E}(\alpha)$ and its non-erasing pattern language $L_{NE}(\alpha)$ are defined respectively by
$L_{E}(\alpha)$ $=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha)\},\text{ and }$
$L_{NE}(\alpha)$ $=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\}.$
@@
Let $RegPat_{\Sigma}=\{\alpha\in Pat_{\Sigma}\mid x_{i}\in\operatorname{alph}(\alpha)\cap X,|\alpha|_{x_{i}}=1\}$ denote the set of all regular patterns, i.e., the set of all patterns where each variable occurs only once. Let $R$ be some set of relations and denote for each $n\in\mathbb{N}$ by $R_{n}\subset R$ the subset of $n$-ary relations in $R$. Define $Rel_{R}=\{(r,x_{i_{1}},...,x_{i_{n}})\mid r\in R_{n},x_{i_{1}},...,x_{i_{n}}\in X,i_{1},...i_{n}\in\mathbb{N},n\in\mathbb{N}\}$ to be the set of all combinations to relate variables in $X$ over the set of relations $R$. Then, a pair $(\alpha,r_{\alpha})\in RegPat_{\Sigma}\times 2^{Rel_{R}}$ is a relational pattern over the terminal alphabet $\Sigma$ and the set of relations $R$. We denote the set of all relational patterns over $\Sigma$ and $R$ by $\mathtt{RelPat}_{\Sigma,R}$. For some $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $h\in H$, we say that $h$ is a $r_{\alpha}$-valid substitution if for all $(r^{\prime},x_{i_{1}},...,x_{i_{n}})\in r_{\alpha}$ we have $(h(x_{i_{1}}),...,h(x_{i_{n}}))\in r^{\prime}$, i.e., all variables are substituted by words related to each other regarding the relation $r^{\prime}$. The set of all $r_{\alpha}$-valid substitutions w.r.t. $\Sigma$ is denoted by $H_{\Sigma,r_{\alpha}}$. If $\Sigma$ or $r_{\alpha}$ are apparent, we may also just write $H_{r_{\alpha}}$ or $H$. For any $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, we denote the erasing relational pattern language $L_{E}(\alpha,r_{\alpha})$ and the non-erasing relational pattern language $L_{NE}(\alpha,r_{\alpha})$ of $(\alpha,r_{\alpha})$ respectively by
$L_{E}(\alpha,r_{\alpha})$ $=\{\ h(\alpha)\ |\ h\in H_{\Sigma,r_{\alpha}}\ \}$
$L_{NE}(\alpha,r_{\alpha})$ $=\{\ h(\alpha)\ |\ h\in H_{\Sigma,r_{\alpha}}\text{ and }h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\ \}.$
@@
In this paper, we primarily investigate the considered decision problems under binary relations, i.e., $|R|=1$ and, if $R=\{r\}$, then $r$ is 2-ary/binary. In particular, we are interested in the following selection of specific customary binary relations. Let $\mathcal{R}$ be the set containing all of the following relations.